OpenStudy (hitaro9):

Show that 1.19^n / n^5 diverges with either the ratio test or the root test.

OpenStudy (kc_kennylau):

That means find \(\Large\lim\limits_{n\rightarrow\infty}\dfrac{1.19^{n+1}(n+1)^5}{1.19^nn^5}\)